8. The economy has total checkable deposits of $800 billion, currency held in public is $200 billion, and the banks’ total reserves are $100 billion, while the required reserve ratio is 10%. The amount of Excess Reserves are:

a.) 10 billion

b.) 100 billion

c.) 80 billion

d.) 20 billion

hint "option (b) is NOT correct"
Answer #1

d.) 20 billion

(Required reserves = Required reserve ratio*Checkable deposits = (10%)*800 = $80 billion
Excess Reserves = Total reserves - Required reserves = 100 - 80 = $20 billion)